An electric kettle containing 600 cm3 of water at 90C and with a heater

coil resistance equal to 16 Ω was left connected to the mains. In how

much time will all the water in the kettle boil away? The voltage in the

mains is 120 V and the efficiency of the kettle 60%.


1
Expert's answer
2021-06-18T11:23:34-0400

"V = 600cm^3"

"\rho= 1\frac{g}{cm^3}"

"m= \rho V=600g=0.6kg"

"\text{The amount of energy required to heat water to }100\degree"

"\Delta T =100\degree-9\degree= 81\degree"

"c=4200\frac{J}{kg*C\degree}"

"Q_1= cm\Delta T"

"Q_1= 4200*0.6*81=204120J"

"\text{The amount of energy required to convert water to steam}"

"L =2300000\frac{J}{kg}"

"Q_2 = Lm"

"Q_2=2300000*0.6=1380000J"


"Q=Q_1+Q_2"

"Q=204120+1380000=1584120J"


"\text {according to Joule-Lenz Law:}"

"Q_3=I^2Rt"

"I =\frac{U}{R}"

"Q_3=\frac{U^2t}{R}"

"\eta=0.6"

"Q =Q_3*\eta"

"t = \frac{QR}{U^2\eta}"


"t=\frac{1584120*12}{120^2*0.6}=2200s"


"\text{Answer: }2200\ s"